# Unveiling Linguistic Regions in Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated considerable cross-lingual alignment and generalization ability. Current research primarily focuses on improving LLMs’ cross-lingual generalization capabilities. However, there is still a lack of research on the intrinsic mechanisms of how LLMs achieve cross-lingual alignment. From the perspective of region partitioning, this paper conducts several investigations on the linguistic competence of LLMs. We discover a core region in LLMs that corresponds to linguistic competence, accounting for approximately 1% of the total model parameters. Removing this core region by setting parameters to zero results in a significant performance decrease across 30 different languages. Furthermore, this core region exhibits significant dimensional dependency, perturbations to even a single parameter on specific dimensions leading to a loss of linguistic competence. Moreover, we discover that distinct regions exist for different monolingual families, and disruption to these specific regions substantially reduces the LLMs’ proficiency in those corresponding languages. Our research also indicates that freezing the core linguistic region during further pre-training can mitigate the issue of catastrophic forgetting (CF), a common occurrence observed during further pre-training of LLMs. Overall, exploring the LLMs’ functional regions provides insights into the foundation of their intelligence.
@@
[…]
@@ 1 Introduction
Over the years, the field of Natural Language Processing (NLP) has been at the forefront of understanding the core principles of intelligence. The emergence of Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023), PaLM 2 (Anil et al., 2023), and LLaMA 2 (Touvron et al., 2023), showcases a significant breakthrough. Thanks to unparalleled scales of model architecture and the vastness of training data, these LLMs now exhibit exceptional linguistic competence and can execute complex tasks requiring abstract knowledge (Dong et al., 2023) and reasoning (Cobbe et al., 2021).
@@
Previous research has revealed that LLMs naturally capture cross-linguistic similarities in their representation space, facilitating zero-shot cross-lingual transfer (Pires et al., 2019; Wu and Dredze, 2019; Xu et al., 2023). The model is fine-tuned on one language, enabling the acquisition of comparable capabilities in another language (Muennighoff et al., 2023; Ye et al., 2023), and exhibits the phenomenon of code-switching when generating context, switching between languages within a single utterance (Khanuja et al., 2020; Zhao et al., 2024). Attempts to improve LLMs’ cross-lingual generalization abilities have been successful through parameter and information transfer learning (Üstün et al., 2020; Choenni et al., 2023), aligning languages compulsorily (Sherborne and Lapata, 2022; Shaham et al., 2024) and utilizing in-context learning techniques (Winata et al., 2021; Tanwar et al., 2023). However, a detailed research of underlying the internal mechanisms of how LLMs possess cross-linguistic alignment capability remains elusive.
@@
To delve deeper into the intrinsic mechanisms of LLMs’ linguistic competence, this paper focus on the LLMs’ parameter importance and investigate the linguistic regions of LLMs based on 30 distinct languages’ performance, with the purpose of figuring out the following questions:
@@
Q1: Does a core linguistic region exist within LLMs that facilitates cross-lingual generalization? Through pre-training across six languages and measurement of parameter importance (Section 2.2), we discover a core region in LLMs corresponding to linguistic competence, which accounts for approximately 1% of the model’s total parameters. As shown at the top of Figure 1, removing this region (setting parameters to zero) consistently lead to a significant decline in performance across 30 test languages (Section 3.2).
@@
Furthermore, by visualizing the core linguistic region (Figure 2), we observe that the linguistic core region of LLMs exhibits significant dimensional dependency. In certain dimensions, only perturbing a single parameter could lead to the model losing its linguistic competence (Section 3.3).
@@
Q2: Beyond the core multilingual region within LLMs , do distinct monolingual (or monolingual family) regions exist that specifically influence individual languages? While LLMs possess strong multilingual capabilities, we discover that each individual language (or language family) encompasses independent regions within the LLMs. As shown in the middle of Figure 1, the analysis of the Russian sentences identifies a particular linguistic region that likewise exerts influence both on the Russian and Ukrainian language, both of which belong to the Slavic group (Section 3.4).
@@
Q3: If and how core linguistic regions affect further pre-training, how to utilize it to optimize further pre-training? After pre-training, core linguistic parameter regions of the LLMs are established for multilingual alignment. Notable shifts in these regions potentially lead to a decline in model lingual capabilities. Our findings reveal that freezing this core region can mitigate the issue of catastrophic forgetting (McCloskey and Cohen, 1989; Kemker et al., 2018), a common occurrence observed during further pre-training of LLMs. As shown at the bottom of Figure 1, we investigate the impact of selectively freezing 5% key parameters of all parameters fine-tuning during further pre-training, compared to the full-scale fine-tuning technique. Findings indicate that this method facilitated comparable or even more efficient learning of the target language, while concurrently decelerating the rate of language attrition for previously learned languages (Section 5). Significantly, our methodology is compatible with the data-replay techniques (Robins, 1995), with no necessity for integrating extra components into the model. Unlike regularization methods (Srivastava et al., 2014; Goodfellow et al., 2014), our approach restricts to a minimal core region in LLMs.
@@
  • We observe that distinct monolingual regions exist in LLMs for different languages (or language families). Removing a specific monolingual region causes a significant deterioration in the linguistic capabilities within that language (or language family).
  • We perform further pre-training for specific languages within the core linguistic region of LLMs frozen, achieving comparable or even superior performance in the target language while mitigating catastrophic forgetting in non-target languages.
@@ 2.1 Model Pre-training
Pre-training is a crucial process by which LLMs acquire linguistic competence and gain general knowledge about the real world. Formally, given a large corpus $\mathcal{D}$, the training objective is to find the optimal $\theta$, minimize the following loss $\mathcal{L}$:
@@ 2.2 Parameter Importance
Calculating $\mathcal{I}_{j}(\theta)$ for each parameter, as outlined in 2, is computationally expensive because it involves $d$ distinct versions of the network computing, for each removed parameter. This becomes particularly challenging as the number of model parameters, $d$, grows to hundreds of billions. However, according to the Taylor expansion formula for $\mathcal{L}$ at $\theta_{j}=0$:
@@ 3.2 Core Linguistic Competence Region
In this section, we conduct further pre-training experiments on LLaMA-2 across six languages, aiming to explore and identify core parameter regions associated with linguistic competence. Specifically, according to Equation 4, we cumulatively compute *$\mathcal{I}^{*}(\theta)=\Sigma\mathcal{I}(\theta)$ values across six different languages’ training, positing that the set of parameters exhibiting maximal importance score *$\mathcal{I}^{*}(\theta)$ during the language further pre-training may have a strong correlation with the model’s linguistic competence, and we provide both logical and empirical evidence to support this hypothesis.
@@
Figure 2: Visualization of the linguistic competence region (the ‘Top’ 5% region). The scale from 0 to 1 (after normalization) represent the proportion of parameters within a $3\times 3$ vicinity that belong to the Top region.
@@
In the experiment corresponding to Table 2, we initially zero out various regions within LLaMA. Consistent with the findings from Table 1, removing the ‘Top’ region leads to a loss of linguistic competence, whereas the ‘Bottom’ region don’t. However, in this experiment, we sought to ascertain if LLaMA could reacquire its lost cross-lingual generalization competence. Thus, we train on different amounts of Chinese Zhihu corpus and evaluate on Chinese Wechat and English Falcon corpora. The results indicate that unlike the ‘Bottom’ region, if the ‘Top’ region is removed and frozen, the model have to relearn basic language rules in other regions based on the provided Chinese Zhihu corpus, but these rules are inherently biased towards Chinese. Consequently, while its proficiency in Chinese is restored, the English perplexity remains high ($5301.527$). If the ‘Top’ region is removed but not frozen, the model can rebuild its linguistic competence in-place. As its proficiency in Chinese is restored, so is its proficiency in English. This implies that the ‘Top’ region encodes generalizable fundamental linguistic competence. When ‘Top’ region is zeroed-out and frozen, other regions significantly adapt to regain core linguistic competence. Similar conclusions can be obtained with an expanded removal ratio 5%. For further details, see Appendix C.
@@
Table 3: Perplexity of LLaMA-2 after removing certain dimensions in the Attention and Feedforward layers. Here, $N_{d}$ denotes the number of dimensions to remove, ’Top’, ’Middle’, and ’Bottom’ refer to the dimensions with the most, moderate, and least cumulated $\mathcal{I}_{\theta}$ during further pre-training. ’Random’ denotes an equivalent number of dimensions chosen randomly for comparison.
@@ 3.3 Dimensional Dependence
To provide a more intuitive revelation of the spatial distribution characteristics of the linguistic competence region within the model, we visualize the ‘Top’ region. As shown in Figure 2, whether in the attention mechanism layer or the feed-forward layer, the linguistic region displays a distinct concentration in both the rows and columns of the matrices. Such distribution features seem to imply that the model’s linguistic competence is concentrated in specific dimensions.
@@
Instead of discretely removing different unstructured parameters, we selectively remove structured certain rows or columns for each matrix, especially those dimensions encompassing a significant number of ‘Top’ region parameters, termed as ‘Top’ dimensions. As illustrated in Table 3, we attempt to remove the columns of FFN.down and Attn.k/q/v, as well as the rows of Attn.o. The results indicate that removing just these ‘Top’ dimensions leads to a substantial decline in the model’s linguistic competence. However, disturbances to the ‘Middle’ ,‘Bottom’ and ‘Random’ dimensions do not yield noticeable effects. Selecting the dimensional region only from the Attention matrix or inverting rows and columns removals lead to similar findings, as described in D.
@@
Here, we explore whether a specific dimension significantly impact the model’s linguistic competence. As illustrated in Figure 3, we iterate through the key dimensions mentioned in Section 3.3, attempting to perturb the same dimension (random initialization) across all Transformer layers. The results indicate that the impact of dimensions 2100 and 4743 on the LLaMA-2-13B substantially surpassed other dimensions, even when compared to the other three in the ‘Top5’ dimensions. In contrast, perturbing two randomly selected dimensions, 2800 and 4200, yield linguistic performance almost indistinguishable from the unperturbed state.
Figure 3: Perplexity of the LLaMA-2-13B when perturbing the same single dimension (Att.O and FFN.down matrices) across all layers. ‘Topk’ represents the top k dimensions that disrupt the model the most. ‘Random selected’ refers to a randomly chosen dimension. ‘Original’ indicates that no dimensions are disrupted.
@@
We discover that even a slight modification to a single parameter in a model with over 13 billion parameters can lead to a significant decline in its output quality. In Table 4, merely resetting the $2100$-th parameter in the ‘Input_LayerNorm’ module of the $1$-st layer to its initial value causes LLaMA-2-13B’s PPL value to skyrocket from 5.865 to 83224.078. If this weight parameter is multiplied by $10$, the PPL value also rises to $4363.462$. This suggests that even minor changes to a single parameter can cause the model to lose nearly all of its linguistic competence. However, randomly altering the parameters at dimensions 2800 and 4200 doesn’t noticeably impact the model. For more details, refer to Appendix E.
@@
To visually illustrate the impact of the linguistic competence region on the model’s output quality, we use "Fudan University is located in" as a premise and observe the model’s outputs under different parameter perturbations. The results are shown in Figure 4. Compared to randomly selected dimension 4200-th, perturbing model on 2100-th dimension significantly leads to model loses its linguistic competence, producing error or even nonsensical strings.
3.4 Monolingual Family Region
In this section, we wonder if LLMs possess distinct regions within an individual language (or language family). Unlike the core linguistic regions described in Section 3.2, a monolingual family region only has a strong correlation with certain languages, and removing it will only cause significant influence on LLMs’ proficiency in those corresponding languages.
Perturbation Parameter Perplexity
- - 5.865
Reset 1 L1-N2100 83224.078
Reset 1 L1-N2800 5.860
Reset 1 L1-N4200 5.858
Mul 10 L1-N2100 4363.462
Mul 10 L1-N2800 5.859
Mul 10 L1-N4200 5.864

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated considerable cross-lingual alignment and generalization ability. Current research primarily focuses on improving LLMs’ cross-lingual generalization capabilities. However, there is still a lack of research on the intrinsic mechanisms of how LLMs achieve cross-lingual alignment. From the perspective of region partitioning, this paper conducts several investigations on the linguistic competence of LLMs. We discover a core region in LLMs that corresponds to linguistic competence, accounting for approximately 1% of the total model parameters. Removing this core region by setting parameters to zero results in a significant performance decrease across 30 different languages. Furthermore, this core region exhibits significant dimensional dependency, perturbations to even a single parameter on specific dimensions leading to a loss of linguistic competence. Moreover, we discover that distinct monolingual regions exist for different languages, and disruption to these specific regions substantially reduces the LLMs’ proficiency in those corresponding languages. Our research also indicates that freezing the core linguistic region during further pre-training can mitigate the issue of catastrophic forgetting (CF), a common phenomenon observed during further pre-training of LLMs. Overall, exploring the LLMs’ functional regions provides insights into the foundation of their intelligence.
@@
Zhihao Zhang1thanks:  Equal contributions., Jun Zhao1, Qi Zhang13thanks:  Corresponding authors., Tao Gui2, Xuanjing Huang1 1 School of Computer Science, Fudan University 2 Institute of Modern Languages and Linguistics, Fudan University 3 Shanghai Collaborative Innovation Center of Intelligent Visual Computing {zhangzhihao19, zhaoj19, qz, tgui, xjhuang}@fudan.edu.cn
@@ 1 Introduction
Over the years, the field of Natural Language Processing (NLP) has been at the forefront of understanding the core principles of intelligence. The emergence of Large Language Models (LLMs) such as GPT-4 (OpenAI, 2023), PaLM 2 (Anil et al., 2023) and LLaMA 2 (Touvron et al., 2023), showcase a significant breakthrough. Thanks to unparalleled scales of model architecture and the vastness of training data, these LLMs now exhibit exceptional linguistic competence and can execute complex tasks requiring abstract knowledge (Dong et al., 2023) and reasoning (Cobbe et al., 2021).
@@
Previous research has revealed that LLMs naturally capture cross-linguistic similarities in their representation space, facilitating zero-shot cross-lingual transfer (Pires et al., 2019; Wu and Dredze, 2019; Xu et al., 2023). The model is fine-tuned on one language, enabling the acquisition of comparable capabilities in another language (Muennighoff et al., 2023; Ye et al., 2023), and exhibits the phenomenon of code-switching when generating context (Khanuja et al., 2020; Zhao et al., 2024). Attempts to improve LLMs’ cross-lingual generalization abilities have been successful through parameter and information transfer learning (Üstün et al., 2020; Choenni et al., 2023), aligning languages compulsorily (Sherborne and Lapata, 2022; Shaham et al., 2024) and utilizing in-context learning techniques (Winata et al., 2021; Tanwar et al., 2023). However, a detailed investigation into the internal mechanisms of how LLMs possess cross-linguistic alignment capability remains elusive.
@@
To delve deeper into the intrinsic mechanisms of LLMs’ linguistic competence, this paper focuses on the LLMs’ parameter importance and investigate the linguistic regions of LLMs based on 30 distinct languages’ performance, with the purpose of figuring out the following questions:
@@
Q1: Does a core linguistic region exist within LLMs that facilitates cross-lingual generalization? By conducting further pre-training across six languages and evaluating models’ parameter importance (Section 2.2), we discover a core region in LLMs corresponding to linguistic competence, which accounts for approximately 1% of the model’s total parameters. As shown at the top of Figure 1, removing this region (setting parameters to zero) consistently leads to a significant decline in performance across 30 test languages (Section 3.2).
@@
Furthermore, by visualizing the core linguistic region (Figure 2), we observe that the linguistic core region of LLMs exhibits significant dimensional dependency. In certain dimensions, only perturbing a single parameter could lead to the model losing its linguistic competence (Section 3.3). Additionally, ablation study in 3.3 shows that beyond outlier dimensions, other non-outlier dimensions in this region are also critical.
@@
Q2: Beyond the core linguistic region within LLMs , do distinct monolingual regions exist that specifically influence individual languages? While LLMs possess strong multilingual capabilities, we discover that each individual language (or language with similar compositional elements or grammatical structures) encompasses independent regions within the LLMs. As shown in the middle of Figure 1, the analysis of the Russian sentences identifies a particular linguistic region that likewise exerts influence both on the Russian and Ukrainian language, both of which belong to the Slavic group (Section 3.4).
@@
Q3: If and how core linguistic regions affect further pre-training, how to utilize it to optimize further pre-training? After pre-training, core linguistic parameter regions of the LLMs are established for multilingual alignment. Notable shifts in these regions potentially lead to a decline in model lingual capabilities. Our findings reveal that freezing this core region can mitigate the issue of catastrophic forgetting (McCloskey and Cohen, 1989; Kemker et al., 2018), a common phenomenon observed during further pre-training of LLMs. As shown at the bottom of Figure 1, we investigate the impact of selectively freezing 5% key parameters of all parameters during further pre-training, compared to the full-scale fine-tuning technique. Findings indicate that this method facilitates comparable learning of the target language while concurrently decelerating the rate of language attrition for previously learned languages (Section 3.5). Significantly, our methodology is compatible with the data-replay techniques (Robins, 1995; Wei et al., 2023), with no necessity for integrating extra components into the model. Unlike regularization methods (Srivastava et al., 2014; Goodfellow et al., 2014), our approach restricts to a minimal core region in LLMs.
@@
  • We observe that distinct monolingual regions exist in LLMs for different languages. Removing a specific monolingual region causes a significant deterioration in the linguistic capabilities within corresponding language.
  • We perform further pre-training for specific languages within the core linguistic region of LLMs frozen, achieving comparable performance in the target language while mitigating catastrophic forgetting in non-target languages.
@@ 2.1 Model Pre-training
Pre-training is a crucial process by which LLMs acquire linguistic competence and gain general knowledge about the real world. Formally, given a large corpus $\mathcal{D}$, the training objective for auto-regressive language modeling is to find the optimal $\theta$ that minimizes the following loss $\mathcal{L}$:
@@ 2.2 Parameter Importance
Calculating $\mathcal{I}_{j}(\theta)$ for each parameter, as outlined in 2, is computationally expensive because it involves $d$ distinct versions of the network computing, for each removed parameter. This becomes particularly challenging as the number of model parameters, $d$, grows to hundreds of billions. However, similar to several prior works (Molchanov et al., 2019; Zhang et al., 2022), using the Taylor expansion formula for $\mathcal{L}$ at $\theta_{j}=0$:
@@ 3.2 Core Linguistic Competence Region
In this section, we conduct further pre-training experiments on LLaMA-2 across six languages, aiming to explore and identify the core linguistic region in LLMs. Here, we define it as "core linguistic", attributed to its minimal proportion within LLMs, constituting only 1% of the total parameters, and its association with model’s linguistic competence modeling.
Specifically, according to Equation 4, we cumulatively compute $\mathcal{I}^{*}(\theta)=\Sigma\mathcal{I}(\theta)$ values across six different languages’ training, positing that the set of parameters exhibiting maximal importance score $\mathcal{I}^{*}(\theta)$ during the language further pre-training may have a strong correlation with the model’s linguistic competence. We provide both logical and empirical evidence to support this hypothesis.
@@
Figure 2: Visualization of the linguistic competence region (the ‘Top’ 5% region). The scale from 0 to 1 (after normalization) represent the proportion of parameters within a $3\times 3$ vicinity that belong to the ‘Top’ region.
@@
In the experiment corresponding to Table 2, we initially zero out various regions within LLaMA. Consistent with the findings from Table 1, removing the ‘Top’ region leads to a loss of linguistic competence, whereas the ‘Bottom’ region don’t. However, in this experiment, we sought to ascertain if LLaMA could reacquire its lost cross-lingual generalization competence. Thus, we train on different amounts of Chinese Zhihu corpus and evaluate on Chinese Wechat and English Falcon corpora. The results indicate that unlike the ‘Bottom’ region, if the ‘Top’ region is removed and frozen, the model have to relearn basic language rules in other regions based on the provided Chinese Zhihu corpus, but these rules are inherently biased towards Chinese. Consequently, while its proficiency in Chinese is restored, the English perplexity remains high ($5301.527$). If the ‘Top’ region is removed but not frozen, the model can rebuild its linguistic competence in-place. As its proficiency in Chinese is restored, so is its proficiency in English. This implies that the ‘Top’ region encodes generalizable fundamental linguistic competence. When ‘Top’ region is zeroed-out and frozen, other regions difficultly adapt to regain core linguistic competence.
@@
Table 3: Perplexity of LLaMA-2 after removing certain dimensions in the Attention and Feedforward layers. Here, $N_{d}$ denotes the number of dimensions to remove, ‘Top’, ‘Middle’, and ‘Bottom’ refer to the dimensions with the most, moderate, and least cumulated $\mathcal{I}_{\theta}$ during further pre-training. ‘Random’ denotes an equivalent number of dimensions chosen randomly for comparison.
@@ 3.3 Dimensional Dependence
To provide a more intuitive revelation of the spatial distribution characteristics of the linguistic competence region within the model, we visualize the ‘Top’ region. As shown in Figure 2, whether in the attention mechanism layer or the feed-forward layer, the linguistic region displays a distinct concentration in both the rows and columns of the matrices. In Appendix B, we also discover that in various layers, the core linguistic region is concentrated on different heads of the Attn.o matrix. Such distribution features seem to imply that the model’s linguistic competence is localized in specific rows and columns.
@@
Instead of discretely removing different unstructured parameters, we selectively remove structured certain rows or columns for each matrix, especially those dimensions encompassing a significant number of ‘Top’ region parameters, termed as ‘Top’ dimensions. As illustrated in Table 3, we attempt to remove the columns of FFN.down and Attn.k/q/v, as well as the rows of Attn.o. The results indicate that removing just these ‘Top’ dimensions leads to a substantial decline in the model’s linguistic competence. However, disturbances to the ‘Middle’ ,‘Bottom’ and ‘Random’ dimensions do not yield noticeable effects. Selecting the dimensional region only from the Attention matrix or inverting rows and columns removals lead to similar findings, as described in C.
@@
Here, we explore whether a specific dimension significantly impacts the model’s linguistic competence. As illustrated in Figure 3, we iterate through the key dimensions mentioned in Section 3.3, attempting to perturb (random initialization) the same dimension across all Transformer layers. The results indicate that the impact of dimensions 2100 and 4743 on the LLaMA-2-13B substantially surpasses other dimensions, even when compared to the other three in the ‘Top5’ dimensions. In contrast, perturbing two randomly selected dimensions, 2800 and 4200, yields linguistic performance almost indistinguishable from the unperturbed state.
Figure 3: Perplexity of the LLaMA-2-13B when perturbing the same single dimension (Att.O and FFN.down columns) across all layers. ‘Top k’ represents the top k dimensions that disrupt the model the most. ‘Random selected’ refers to a randomly chosen dimension. ‘Original’ indicates that no dimensions are disrupted.
@@
We discover that even a slight modification to a single parameter in a model with over 13 billion parameters can lead to a significant decline in its output quality. In Table 8, merely resetting the $2100$-th parameter in the ‘Input_LayerNorm’ module of the $1$-st layer to its initial value causes LLaMA-2-13B’s PPL value to skyrocket from 5.865 to 83224.078. If this weight parameter is multiplied by $10$, the PPL value also rises to $4363.462$. However, randomly altering the parameters at dimensions 2800 and 4200 doesn’t noticeably impact the model. For more details, refer to Appendix D.
@@
To visually illustrate the impact of the linguistic competence region on the model’s output quality, we use "Fudan University is located in" as a premise and observe the model’s outputs under different parameter perturbations. The results are shown in Figure 4. Compared to randomly selected 4200-th dimension, perturbing model on 2100-th dimension significantly leads to model loses its linguistic competence, producing error or even nonsensical strings.
Ablation study
Considering that the collapse of PPL may be possibly caused solely by the removal of outlier dimension parameters, rather than the collective effect of the entire linguistic region, we conduct an ablation experiment without removing the outlier dimensions’ parameters (1512/2533 for LLaMA-2-7B and 2100/4743 for LLaMA-2-13B). The results for the 13B model, presented in Table 4, indicate that while outlier dimensions are contained within the core linguistic region, non-outlier dimensions in this region are also critical. Not altering any row or column of the matrix parameters for the two outlier dimensions also results in a great increase in PPL, although the collapse of PPL is slowed. This finding reveals that all dimensions in the core linguistic region are tightly interrelated, and disrupting even a small part of it can lead to a PPL collapse. For a more detailed analysis of the experiments and results for the 7B model, please refer to Appendix E.
Removal Region $N_{d}$ LLaMA-2-13B Top(100K)
w/ outlier $d$ w/o outlier $d$
Attn.o(row) Attn.k/q/v(column) FFN.down(coulumn) 1 5218.1 11.079
3 37344.078 77.519
5 41840.613 590.895
10 62579.516 513998.437
Attn.o(row) Attn.k/q/v(column) 1 10.899 10.901
3 44.384 44.389
5 33.52 29.793
10 118.968 120.977
Attn.o(column) Attn.k/q/v(row) 1 17.609 13.666
3 313.178 63.99
5 526.464 163.388
10 5841.446 2675.347
FFN.up/gate(row) FFN.down(column) 1 154.925 6.142
3 33995.949 6.668
5 32572.888 8.139
10 524867.687 45.408
Table 4: Perplexity of LLaMA-2-13B after removing ‘Top’ certain dimensions w/ or w/o outlier dimensions respectively. Here, $N_{d}$ denotes the number of dimensions to remove, ‘Top’ refers to the dimensions with the most cumulated $\mathcal{I}_{\theta}$ during further pre-training.
3.4 Monolingual Region
In this section, we wonder if LLMs possess distinct regions within an individual language. Unlike the core linguistic regions described in Section 3.2, a monolingual family region only has a strong correlation with certain languages, and removing it will only cause significant influence on LLMs’ proficiency in those corresponding languages.